OpenStudy (anonymous):

Please help!! I will give medals!! I need help catching up Question 2. 2. If a sample of nitrogen gas (N2) contains 3.0 × 1024 atoms of nitrogen, then about how many moles of nitrogen gas are in the sample? (Points : 3) 1.0 2.5 3.0 5.0

OpenStudy (anonymous):

@wio @sammixboo @camerondoherty

OpenStudy (displayerror):

Use Avogadro's number: 6.02 x 10^23 of any item is in one mole of that item. Thus, one mole of potatoes contains 6.02 x 10^23 potatoes. Using that same reasoning, 1 mole of Nitrogen would contain 6.02 x 10^23 atoms of Nitrogen \[3.0 \times 10^{24} \ \text{atoms of Nitrogen} \times \frac{1 \ \text{mole of Nitrogen}}{6.02 \times 10^{23} \ \text{atoms of Nitrogen}}\] In the above equation, when we multiply, the unit "atoms of Nitrogen" cancels out because there's one in the numerator and one in the denominator, leaving us with the desired unit, "moles of Nitrogen".